### v1

# Active Learning for Machine Learning Driven Molecular Dynamics

Abstract
Machine-learned coarse-grained (CG) potentials are fast, but degrade over time when simulations reach under-sampled bio-molecular conformations, and generating widespread all-atom (AA) data to combat this is computationally infeasible. We propose a novel active learning framework for CG neural network potentials in molecular dynamics (MD). Building on the CGSchNet model [7], our method employs root mean squared deviation (RMSD)-based frame selection from MD simulations in order to generate data on-the-fly by querying an oracle during the training of a neural network potential. This framework preserves CG-level efficiency while correcting the model at precise, RMSD-identified coverage gaps. By training CGSchNet—a coarse-grained neural network potential—we empirically show that our framework explores previously unseen configurations and trains the model on unexplored regions of conformational space. Our active learning framework enables a CGSchNet model trained on the Chignolin protein to achieve a 33.05% improvement in the Wasserstein-1 (W1) metric in Time-lagged Independent Component Analysis (TICA) [5] space on an in-house benchmark suite.
1Baskin Engineering, University of California - Santa Cruz, Santa Cruz, CA
2GiwoTech Inc, Cambridge, MA
1 Introduction
Advancements in the realm of MD have allowed for the effective exploration of complex conformational landscapes of biomolecular structures such as proteins, but such methods remain limited by the immense computational overhead associated with system size and the granularity of the atomic representation. [3] Multiple generations of AA neural network potentials have progressed high-resolution MD by taking atomic coordinates, , and predicting the energy with forces , surpassing classical energy functionals in terms of accuracy [1]. However, such high-dimensional models remain slow at AA resolution and are often more computationally expensive than computing classical force-field terms. [3] Consequently, simulating large structures takes a prohibitive amount of time to reach critical and relevant regions, thereby making speedup and efficiency key goals in MD development.
A renowned approach for achieving longer timescales with neural network potentials in MD is to reduce the size of the conformational space through coarse-graining, in which groups of atoms are represented as single beads. Generally, such models have been trained via force matching, which fits the CG potential to match reference AA forces on mapped configurations to minimize the difference between CG forces and mapped AA forces [4]. However, because most training data samples single metastable states in proteins, they struggle to generalize and fully explore the conformational space, particularly during transitions between metastable states, making simulations prone to conformational "explosion" or "implosion" anomalies. These divergences result from the network generating physically inconsistent forces upon encountering configurations significantly different from those in the initial training data. This issue is especially prevalent for large proteins with complex folding structures.
To address this, we propose an active learning framework that uses targeted, on-the-fly data acquisition to obtain data beyond the regions normally sampled by MD by selecting specific out-of-distribution frames during simulation and querying an AA ground-truth oracle, thereby patching coverage gaps at minimal cost. For data acquisition, we identify frames most different from the training set quantified by distance-based proxies such as RMSD, enabling data generation in the least-explored regions of the protein conformational space. We evaluate our approach on the Chignolin protein and observe that active learning achieves an improvement of 33.05% in TICA space compared to a ground truth, and significant qualitative exploration improvement when analyzed with RMSD.
2 Methodology
In Fig. 1 we show an overview of our active learning methodology. At each active learning iteration, the model is trained on MD trajectory data and then used to simulate the CG protein system. We then select the frames with the largest RMSD values, indicating the greatest difference between the configuration and the training dataset, ultimately targeting the least-covered configurations. These frames are then backmapped into AA space, used to query the oracle by seeding OpenMM [2] all-atom simulations, projected back into CG space, and appended to the dataset. The model is retrained, and the loop repeats, querying the oracle only where coverage is low to minimize usage of AA simulations.
2.1 Neural network potential (CGSchNet)
Our neural network leverages a graph neural network (GNN) architecture based on CGSchNet [7], using continuous filter convolutions to build features based on inter-bead distances. Let us define the CG bead coordinates as , and the inter-bead distance as . Based on these distances, the network outputs a scalar energy potential which accounts for the system’s geometry rather than just the absolute position or orientation. This architecture thereby ensures that remains invariant to global translations and rotations, and the resultant forces are translation-invariant and rotation-equivariant as well. The objective of force matching is done by minimizing the mean-squared error between the predicted forces and the target CG forces projected from AA space. The force matching loss function is given as:
where represents the CG bead coordinates for beads in frame , represents the AACG projected CG forces and tells us the number of training frames.
2.2 AA CG and CG AA projection
Our neural network potential is compatible with CG data, while the oracle simulator works only with AA data. Thus, we require a bidirectional bridge between the spaces. We convert AA to CG by mapping atomic coordinates to carbon-alpha () beads via a pre-processing script using a linear operator
We map the forces by projecting AA forces onto CG degrees of freedom via force projection operators [7, 4]:
To convert from CG to AA, we utilize the PULCHRA [6] backmapper to reconstruct the missing, non- atoms into statistically likely positions.
2.3 Frame selection
During each active learning iteration, we compute histograms of RMSD values between simulated frames and the training set, and select the frames with the largest simulation-minus-training RMSD discrepancies as candidates. We then filter further by removing any frames with RMSD value outside of a cutoff to avoid getting frames that are part of implosions or explosions. These frames are then backmapped to AA space in order to seed the OpenMM [2] oracle which generates robust data that is then projected back into CG space to begin retraining.
3 Results
To test the efficacy of our active learning framework, we trained a base model on a small subset of data from the Chignolin protein, and used an in-house benchmark suite to measure the (i) conformation exploration along slow modes of motion via TICA projections [5], (ii) reaction coordinates representing the fraction of time certain atoms are within close proximity, and the distributions of (iii) bond lengths, (iv) bond angles, and (v) dihedral angles. Model-generated distributions were compared against extensive Chignolin ground-truth references, depicted in Figure A.1 for the base model, and in Figure A.2 for the model after applying the active learning loop. We summarize the outcomes quantitatively with the W1 distance metric between distributions, representing the minimal "work" required to transform the model distribution into the reference, where lower values are favorable. We report the W1 metrics before and after applying the active learning loop in Table 1.
| W1 | ||
| Plot | Before AL | After AL |
| TICA KDE | 1.15023 | 0.77003 |
| Reaction coordinate | 0.15141 | 0.38302 |
| Bond length | 0.00043 | 0.00022 |
| Bond angle | 0.11036 | 0.10148 |
| Dihedral | 0.25472 | 0.36378 |
These benchmarks show significant improvements in exploration capability and local accuracy measures. In TICA space, the W1 distance to the reference dropped by 33.05%. Additionally, it produced more stable trajectories, with bond-length and bond-angle distributions aligning more closely with the ground truth with a 48.84% and 8.05% decrease in W1 distance, respectively.
Improvements in exploration are also evident in the RMSD histograms before and after active learning (Figure 2). The base model (black) shows a bimodal RMSD density distribution, indicating sampling concentrated in two conformational states. The additional histograms represent the active learning iterations (purple red), reflecting the data that was used to train the model at each iteration. The substantial variance across iterations indicates that each round, the model targets different under-sampled regions, with the active learning loop supplying training data to support it across diverse conformational states. The post-active learning model (brown) shows a much broader RMSD distribution, further illustrating improvements in phase-space coverage aligning with the TICA W1 reductions reported above.
4 Discussion
Since we query OpenMM only when the RMSD histograms indicate poorly covered regions, we retain the speed and efficiency of CG simulations while correcting the model in those under-sampled states on-the-fly, restoring accuracy and thoroughness in coverage. Evaluated on Chignolin’s complex structure, pure CG simulation without using our active learning framework often resulted in explosions or implosions due to limited coverage and generalization. Empirically, we found that our active learning method explored more of the conformational space and improved accuracy across the bond, bond-angle, dihedral, and TICA space metrics in the standardized benchmark. This indicates that long CG trajectories are stabilized and explosion and implosion anomalies are avoided, and more of the conformation space is explored, which is crucial for the ultimate goal of generalization.
These results demonstrate that targeted active learning in MD stabilizes CG simulations and expands conformational coverage within reasonable compute, while also providing paths for future improvements upon certain mechanisms of our framework. Stronger backmapping methodologies would allow for increased accuracy in AA reconstructions while reducing relaxation cost. Furthermore, more rigorous distance proxies or measures may further optimize how frames are prioritized and, in turn, reduce excessive oracle usage. Ultimately, our active learning framework’s targeted, precise, oracle-backed patching of phase-space gaps can significantly expand the training data for ML models in molecular dynamics at a practical cost, and provides clear paths for improvements in selection, backmapping, and representation bridging.
Impact statement
Accurate yet affordable exploration of protein conformational spaces is a critical bottleneck in ML-driven drug discovery, where rare conformational states and transitions govern crucial protein mechanisms such as binding ability. Our active learning framework unifies high-speed CG neural network potentials with precise all-atom oracle queries to explore conformational spaces more rigorously while keeping the computational cost of labeling low. Our result enables increased capability in viewing protein behavior, which may play a role in revealing unique binding opportunities or promising compounds far sooner, minimizing the time spent in extensive trial-and-error in drug discovery laboratories. Because our framework is model-agnostic and bridges CGAA reliably, it can be used to augment current or future state-of-the-art ML force fields easily rather than replacing them altogether, resulting in confident observations and shorter paths from trial-and-error to effective treatments.
5 Data and Code Availability
The code is available at: github.com/BioMedAI-UCSC/cgschnet/tree/refactor-active-learning
Appendix A Technical Appendices and Supplementary Material